n=10
p=0.48
q=0.52
x=8
how to calculate this:
10c8(0.48)^8(0.52)^2
please let me know.


Answer by jim_thompson5910(35256)

10 C 8 = (10!)/(8!*(10-8)!)

10 C 8 = (10!)/(8!*2!)

10 C 8 = (10*9*8!)/(8!*2!)

10 C 8 = (10*9)/(2!)

10 C 8 = (10*9)/(2*1)

10 C 8 = 45

-------------------------------------------------------

(10 C 8)(0.48)^8(0.52)^2 = (45)*(0.48)^8(0.52)^2

(10 C 8)(0.48)^8(0.52)^2 = (45)*0.0028179280429*0.2704

(10 C 8)(0.48)^8(0.52)^2 = 0.034288548426